# A Game-Theoretic Unital Classification Theorem for $C^{*}$-Algebras

### v1

@@ 1. Introduction
$C^{*}$-classification developed with close relations to these structural results of von Neumann algebras. Indeed, the very first classification result for simple $C^{*}$-algebras is a precise analogue to Murray and von Neumann’s characterization of the hyperfinite II1 factor [murray1943rings], given by Glimm’s classification of uniformly hyperfinite $C^{*}$-algebras [glimm1960]. The modern-day program mirrors Connes’ groundbreaking work classifying amenable von Neumann algebras [connes1976classification], and has been an ongoing endeavor involving many mathematicians since the 1970s. This began with Elliott’s result classifying approximately finite-dimensional (AF) algebras using a relatively simple invariant: their ordered $K_{0}$ groups [elliott1976AF]. The classification program continued with the Kirchberg-Phillips theorem [kirchberg1995classif, phillips2000classif], classifying a large class of algebras by their operator-algebraic $K$-theory. Finally, the state-of-the-art classification result also involves the trace space as a part of the invariant. The following [CGSTW-new-classificationI-2023, Theorem A] is usually attributed to ‘many hands’, and a brief historical overview is given in [CGSTW-new-classificationI-2023, Section 1.2].
@@
We hereafter refer to this class of $C^{*}$-algebras as classifiable algebras. The class directly parallels Connes’ classification of amenable factors; indeed, the assumptions of unitality, simplicity, and separability are immediate in the case of von Neumann factors, and nuclearity is equivalent to amenability in the von Neumann factor setting. The term $\mathcal{Z}$-stable” refers to tensorial absorption of a particular $C^{*}$-algebra called the Jiang-Su algebra [MR1680321], denoted by $\mathcal{Z}$. This assumption is crucial since the $K$-theory of an algebra $A$ and $A\otimes\mathcal{Z}$ are identical. Finally, the assumption of the UCT allows one to connect Kasparov’s $KK$-theory [Kasparov1981-origKK] with $K$-theory, which is necessary to phrase the invariant in terms of $K$-theory. Indeed, the aforementioned Kirchberg-Phillips theorem shows that the purely infinite classifiable algebras are isomorphic if and only if they are $KK$-equivalent. We will return to discuss the UCT assumption shortly after stating Theorem D.
@@
Theorem 1.1 says that two classifiable $C^{*}$-algebras are isomorphic if and only if their invariants are isomorphic. It is sensible to ask whether the classification theorem can be extended to coarser equivalence relations between separable $C^{*}$-algebras, respectively between their invariants. We answer precisely this question in the positive, for a family of equivalence relations arising from the infinitary logic $L_{\omega_{1},\omega}$. When $L$ is a discrete first-order language, the set of $L_{\omega_{1},\omega}$-formulas is obtained by closing the set of first-order (finitary) $L$-formulas under countable conjunctions and disjunctions. The quantifier-rank of an $L_{\omega_{1},\omega}$-formula $\varphi$ is a countable ordinal $\operatorname{qr}(\varphi)<\omega_{1}$ corresponding roughly to the maximal depth of nested quantifier-chains in $\varphi$. Setting $A\equiv_{\alpha}B$ whenever $A$ and $B$ are $L$-structures that agree on $L_{\omega_{1},\omega}$-sentences of quantifier rank at most $\alpha$ yields a sequence $(\equiv_{\alpha})_{\alpha<\omega_{1}}$ of refining equivalence relations. Scott’s isomorphism theorem [Scott1965-SCOLWD] states that the intersection of this sequence coincides with isomorphism, that is, countable $L$-structures AA and $B$ are isomorphic (written $A\cong B$) if and only if $A\equiv_{\alpha}B$ for every $\alpha<\omega_{1}$. To study metric structures such as $C^{*}$-algebras in the context of (infinitary) logic, we use metric model theory as developed by Ben Yaacov, Berenstein, Henson and Usvyatsov [MR2436146]. The corresponding generalization of $L_{\omega_{1},\omega}$ due to Ben Yaacov, Doucha, Nies and Tsankov [yaacov2017metric] serves as the basic framework of this paper, providing in particular a metric version of the equivalence relations $(\equiv_{\alpha})_{\alpha<\omega_{1}}$ and of Scott’s isomorphism theorem.
@@
In [games_on_AF_algebras], the authors prove the following generalisation of Elliott’s classification of AF-algebras, in terms of infinitary metric logic.
Theorem 1.2 ([games_on_AF_algebras, Theorem A]).
@@
The proof from [games_on_AF_algebras] combines the intertwining argument of the Elliott classification program with a game-theoretic description of $\equiv_{\alpha}$. This naturally leads to the question of whether the same can be done for unital Kirchberg algebras satisfying the UCT, see [games_on_AF_algebras, Question 1.2]. In this paper, we not only answer this question, but also extend the result to all classifiable $C^{*}$-algebras, and establish the converse, answering in particular [games_on_AF_algebras, Question 4.5]. Below, we write $KT_{u}(\cdot)$ for the functor taking a $C^{*}$-algebra to its classifying invariant consisting of $K$-theory and traces. We lay this out more explicitly in Section 3.2, where we also explain how one can naturally define relations $\equiv_{\alpha}$ between classifying invariants of $C^{*}$-algebras. Our main result is the following.
@@
There exists a function $\theta\colon\omega_{1}\rightarrow\omega_{1}$ such that for every pair of classifiable $C^{*}$-algebras $A$ and $B$ and every $\alpha<\omega_{1}$,
$KT_{u}(A)\equiv_{\theta(\alpha)}KT_{u}(B)\quad\implies\quad A\equiv_{\alpha}B.$
Furthermore, there exists a function $\theta^{\prime}\colon\omega_{1}\rightarrow\omega_{1}$ such that for every $\alpha<\omega_{1}$,
@@ 1.1. Methods
Rather than going directly through the intertwining arguments of $C^{*}$-classification, our approach reduces the proof of Theorem A to checking that the set of classifiable $C^{*}$-algebras, the functor $KT_{u}(\cdot)$, and the relations $\equiv_{\alpha}$ satisfy certain descriptive ‘tameness’ conditions. This allows us to handle all classifiable $C^{*}$-algebras at once, using Theorem 1.1. More precisely, the following theorem proved in Section 2.4 is the descriptive set theoretic heart of our method.
@@
Let $X,Y$ be analytic [^1] An analytic set is the image of a Borel set under a Borel map, and can be thought of as “one level more complex than Borel”. subsets of a standard Borel space and let $K\colon X\to Y$ be an analytic map. Assume that $E\subseteq X^{2}$ and $F\subseteq Y^{2}$ are Borel equivalence relations, and $(E_{\alpha})_{\alpha<\omega_{1}}$ and $(F_{\alpha})_{\alpha<\omega_{1}}$ are refining uniformly analytic equivalence relations on $X$ and $Y$ respectively, and denote $E_{\infty}=\bigcap_{\alpha<\omega_{1}}E_{\alpha},F_{\infty}=\bigcap_{\alpha<\omega_{1}}F_{\alpha}$. Suppose
@@
The term ‘refining’ here means that $E_{\alpha}\supseteq E_{\beta}$ whenever $\alpha<\beta$, and ‘uniformly analytic’ means the equivalence relations extend to a family of analytic equivalence relations indexed by the linear orders (see Definition 2.25). Our proof is non-constructive, and no bound on $\alpha$ and $\beta$ can be extracted in general. For our applications to classification of $C^{*}$-algebras, we derive the following corollary of Theorem B and the metric Scott’s isomorphism theorem from [yaacov2017metric]. Here $L_{i}$ is a metric language and $\operatorname{Mod}_{\omega}(L_{i})$ denotes a set of countable encodings for separable $L_{i}$-structures, for $i=1,2$. The set $\operatorname{Mod}_{\omega}(L_{i})$ carries a natural Borel structure, described in Section 2.5. In the formulation below we identify a code for a separable $L_{i}$-structure $A\in\operatorname{Mod}_{\omega}(L_{i})$ with the structure it encodes.
@@
Let $L$ be a countable metric language. Then the refining family of equivalence relations $(\equiv_{\alpha}^{bf})_{\alpha<\omega_{1}}$ on $\operatorname{Mod}_{\omega}(L)$ is uniformly Borel.
@@
To conclude Theorem A, we are left with showing that the previous corollary can be applied to the set of classifiable $C^{*}$-algebras and the invariant $KT_{u}(\cdot)$. The descriptive set theory of the Elliott invariant, and of various assumptions of the unital classification theorem have been previously studied in [farah2013descriptive]. In particular, using the results of Farah, Toms, and Törnquist, we reduce the proof of Theorem A to establishing that the set of $C^{*}$-algebras satisfying the UCT is an analytic set (with respect to a natural Borel structure on the set of unital separable $C^{*}$-algebras, see Section 3.1).
@@
Rosenberg and Schochet established their universal coefficient theorem (UCT) [RosenbergSchochetUCT], which computes $KK$-theory [^3] For $C^{*}$-algebras $A,B$, $KK(A,B)$ is an abelian group that can be defined in terms of Kasparov $(A,B)$-bimodules. We use a different definition outlined in what follows. in terms of $K$-theory, for a large class of $C^{*}$-algebras. They show that a $C^{*}$-algebra satisfies the UCT if and only if it is $KK$-equivalent to an abelian $C^{*}$-algebra (for our purposes this can be taken as the definition of the UCT). This condition plays a central and mysterious role within the classification of $C^{*}$-algebras; it is unknown whether it is redundant and begs the open question of whether every simple separable nuclear $C^{*}$-algebra satisfies the UCT. One can think of $KK$-equivalence as a loose notion of homotopy equivalence between $C^{*}$-algebras, and we study the descriptive complexity of the UCT class by using the picture of $KK$ as a ‘universal functor’. More precisely, from the work [Higson-KKtheory] of Higson, it follows two $C^{*}$-algebras are $KK$-equivalent if and only if their images under any functor that satisfies certain conditions (that we describe in Section 3.1, see Definition 3.7) become equivalent. The required conditions are universal (i.e., of the form “for all …”), which in turn allows us to present $KK$-equivalence of $C^{*}$-algebras $A,B$ in an existential way: there exist a finite number of conditions on functors $F$, that imply equivalence of images $F(A)\cong F(B)$.
@@
In order to show that $KK$-equivalence is an analytic relation (which implies analyticity of the UCT), we define a first-order language $L_{KK}$ and an $L_{KK}$-theory $T_{KK}$, whose models correspond to functors satisfying Higson’s conditions. We prove by a thorough analysis of Higson’s conditions that $T_{KK}$ is analytic as a subset of the set of $L_{KK}$-sentences with a suitable Borel space structure. This is carried out in Section 3.3 and uses a particular construction of the Borel category $\mathbf{C^{*}}$ of separable $C^{*}$-algebras. More precisely, in Section 2.5 we describe a Borel category equivalent to the category of separable models of a fixed metric $L_{\omega_{1},\omega}$-sentence with $L$-homomorphisms as morphisms, that we use in the case of $C^{*}$-algebras (see Proposition 2.39). Let us note that if one considers only surjective isometries between models as morphisms, such Borel groupoids were already considered in [ElliottEtAl2013, MR3660238] (using the Urysohn space) and in [yaacov2017metric]. As a corollary of analyticity of $T_{KK}$, we get the following, which is a result of interest in its own right.
@@
    It follows from previous works on the descriptive set theory of $C^{*}$-algebra invariants (together with Theorem E), that the set of classifiable $C^{*}$-algebras is analytic, and that the map $KT_{u}(\cdot)$ taking a $C^{*}$-algebra to its invariant is Borel.
@@ Acknowledgements
We are indebted to copious amounts of guidance from Udi Hrushovski and Stuart White. We also thank the participants of the reading group on [games_on_AF_algebras] in Oxford, Michaelmas 2024, in particular Jakub Curda, Shanshan Hua and Austin Shiner. Moreover, we thank Filippo Calderoni and Andrea Vaccaro for some conversations. JP was supported by the Engineering and Physical Sciences Research Council (EP/X026647/1). MS was supported by the National Science Foundation under Grant No. DMS-2424441, and by the IAS School of Mathematics.
@@ 2. Descriptive set theory and infinitary metric logic
The model theory of metric structures, dating back to work of Chang and Keisler [chang_keisler_continuous], was developed extensively in the past two decades, and large parts of (discrete) model theory have been generalized to the metric context [MR2436146, Ben_Yaacov_2010]. Metric model theory has since received growing attention, particularly within the study of operator algebras [farah2021MTofC*, Goldbring2023book, goldbring2022existentially, ioana2024existential, robert2025selfless]. The study of infinitary logics emerged in the 1960s, with particular focus on the logic $L_{\omega_{1},\omega}$, where one allows disjunction and conjunction over countable sets of formulas. [^4] The subscript $\omega_{1}$ indicates that only countable disjunctions and conjunctions appear in $L_{\omega_{1},\omega}$-formulas, while the subscript $\omega$ indicates that only finite-length blocks of quantifiers appear.
We aim to give in Section 2.2 a brief introduction to the lesser studied area of infinitary metric logic as treated in [yaacov2017metric], as well as its connection to descriptive set theory through considering spaces of separable models. We define the notions of a metric language $L$, the corresponding infinitary logic $L_{\omega_{1},\omega}$ and the quantifier rank $\operatorname{qr}(\varphi)$ of an $L_{\omega_{1},\omega}$-formula, and describe the standard Borel space $\operatorname{Mod}_{\omega}(L)$ of separable $L$-structures. We will assume familiarity with some basic concepts in discrete first-order logic.
In section 2.3 we give a presentation of the back-and-forth pseudo-distance of Ben Yaacov, Doucha, Nies, and Tsankov [yaacov2017metric], in terms of a game which we refer to here as the dynamic metric Ehrenfeucht-Fraïssé game $\operatorname{EFD}_{\alpha}$. We show that the set of winning strategies for such games is Borel, and derive in particular that the family of equivalence relations $(\equiv_{\alpha}^{bf})_{\alpha<\omega_{1}}$ between separable structures is Borel in a uniform sense, see Theorem 2.21.

### v2

@@ 1. Introduction
$C^{*}$-classification developed with close relations to these structural results of von Neumann algebras. Indeed, the very first classification result for simple $C^{*}$-algebras is a precise analogue to Murray and von Neumann’s characterization of the hyperfinite II1 factor [46], given by Glimm’s classification of uniformly hyperfinite $C^{*}$-algebras [20]. The modern-day program mirrors Connes’ groundbreaking work classifying amenable von Neumann algebras [8], and has been an ongoing endeavor involving many mathematicians since the 1970s. This began with Elliott’s result classifying approximately finite-dimensional (AF) algebras using a relatively simple invariant: their ordered $K_{0}$ groups [13]. The classification program continued with the Kirchberg-Phillips theorem [42, 47], classifying a large class of algebras by their operator-algebraic $K$-theory. Finally, the general version of the unital classification theorem combines the Kirchberg-Phillips theorem with a stably finite classification theorem which was proved combining [11, 23, 24, 52], and the many papers these rely on. We use the version from the new abstract approach to the stably finite classification [5, Theorem A].
@@
We hereafter refer to this class of $C^{*}$-algebras as classifiable algebras. The class directly parallels Connes’ classification of amenable factors; indeed, the assumptions of unitality, simplicity, and separability are immediate in the case of von Neumann factors, and nuclearity is equivalent to amenability in the von Neumann factor setting. The term $\mathcal{Z}$-stable” refers to tensorial absorption of a particular $C^{*}$-algebra called the Jiang-Su algebra [36], denoted by $\mathcal{Z}$. This assumption is crucial since the $K$-theory of an algebra $A$ and $A\otimes\mathcal{Z}$ are identical. Finally, the assumption of the UCT allows one to connect Kasparov’s $KK$-theory [38] with $K$-theory, which is necessary to phrase the invariant in terms of $K$-theory. Indeed, the aforementioned Kirchberg-Phillips theorem shows that the purely infinite classifiable algebras are isomorphic if and only if they are $KK$-equivalent. We will return to discuss the UCT assumption shortly after stating Theorem D.
@@
Theorem 1.1 says that two classifiable $C^{*}$-algebras are isomorphic if and only if their invariants are isomorphic. It is sensible to ask whether the classification theorem can be extended to coarser equivalence relations between classifiable $C^{*}$-algebras, respectively between their invariants. We answer precisely this question in the positive, for a family of equivalence relations arising from the infinitary logic $L_{\omega_{1},\omega}$. When $L$ is a discrete first-order language, the set of $L_{\omega_{1},\omega}$-formulas is obtained by closing the set of first-order $L$-formulas under countable conjunctions and disjunctions. The quantifier-rank of an $L_{\omega_{1},\omega}$-formula $\varphi$ is a countable ordinal $\operatorname{qr}(\varphi)<\omega_{1}$ corresponding roughly to the maximal depth of nested quantifier-chains in $\varphi$. Setting $A\equiv_{\alpha}B$ whenever $A$ and $B$ are $L$-structures that agree on $L_{\omega_{1},\omega}$-sentences of quantifier rank at most $\alpha$ yields a sequence $(\equiv_{\alpha})_{\alpha<\omega_{1}}$ of refining equivalence relations. Scott’s isomorphism theorem [50] states that the intersection of this sequence coincides with isomorphism, that is, countable $L$-structures AA and $B$ are isomorphic (written $A\cong B$) if and only if $A\equiv_{\alpha}B$ for every $\alpha<\omega_{1}$. To study metric structures such as $C^{*}$-algebras in the context of (infinitary) logic, we use metric model theory as developed by Ben Yaacov, Berenstein, Henson and Usvyatsov [1]. The corresponding generalization of $L_{\omega_{1},\omega}$ due to Ben Yaacov, Doucha, Nies and Tsankov [2] serves as the basic framework of this paper, providing in particular a metric version of the equivalence relations $(\equiv_{\alpha})_{\alpha<\omega_{1}}$ and of Scott’s isomorphism theorem.
@@
In [9], the authors prove the following generalisation of Elliott’s classification of AF-algebras, in terms of infinitary metric logic.
Theorem 1.2 ([9, Theorem A]).
@@
The proof from [9] combines the intertwining argument of the Elliott classification program with a game-theoretic description of $\equiv_{\alpha}$. This naturally leads to the question of whether the same can be done for unital Kirchberg algebras satisfying the UCT, see [9, Question 1.2]. In this paper, we not only answer this question, but also extend the result to all classifiable $C^{*}$-algebras, and establish the converse, answering in particular [9, Question 4.5]. Below, we write $KT_{u}(\cdot)$ for the functor taking a $C^{*}$-algebra to its classifying invariant consisting of $K$-theory and traces. We lay this out more explicitly in Section 3.2, where we also describe a metric logic language for classifying invariants of $C^{*}$-algebras. Our main result is the following.
@@
There exist functions $\theta\colon\omega_{1}\rightarrow\omega_{1}$ and $\theta^{\prime}\colon\omega_{1}\rightarrow\omega_{1}$ such that for every pair of classifiable $C^{*}$-algebras $A$ and $B$ and every $\alpha<\omega_{1}$,
$KT_{u}(A)\equiv_{\theta(\alpha)}KT_{u}(B)\quad$ $\implies\quad$ $A\equiv_{\alpha}B,$
@@ 1.1. Methods
In contrast to [9], rather than going directly through the intertwining arguments of $C^{*}$-classification, our approach reduces the proof of Theorem A to checking that the set of classifiable $C^{*}$-algebras, the functor $KT_{u}(\cdot)$, and the relations $\equiv_{\alpha}$ satisfy certain descriptive ‘tameness’ conditions. This allows us to handle all classifiable $C^{*}$-algebras at once, using Theorem 1.1. More precisely, the following theorem proved in Section 2.4 is the descriptive set theoretic heart of our method.
@@
Let $X,Y$ be analytic [^1] An analytic set is the image of a Borel set under a Borel map. subsets of a standard Borel space and let $K\colon X\to Y$ be an analytic map. Assume that $E\subseteq X^{2}$ and $F\subseteq Y^{2}$ are Borel equivalence relations, and $(E_{\alpha})_{\alpha<\omega_{1}}$ and $(F_{\alpha})_{\alpha<\omega_{1}}$ are refining uniformly analytic equivalence relations on $X$ and $Y$ respectively, and denote $E_{\infty}=\bigcap_{\alpha<\omega_{1}}E_{\alpha},F_{\infty}=\bigcap_{\alpha<\omega_{1}}F_{\alpha}$. Suppose
@@
The term ‘refining’ here means that $E_{\alpha}\supseteq E_{\beta}$ whenever $\alpha<\beta$, and ‘uniformly analytic’ means the equivalence relations extend to a family of analytic equivalence relations indexed by the set of countable linear orders (see Definition 2.25). Our proof is non-constructive, and no bound on $\alpha$ and $\beta$ can be extracted in general. For our applications to classification of $C^{*}$-algebras, we derive the following corollary of Theorem B and the metric Scott’s isomorphism theorem from [2]. Here $L_{i}$ is a metric language and $\operatorname{Mod}_{\omega}(L_{i})$ denotes a suitable encoding of the set of isomorphism classes of separable $L_{i}$-structures, for $i=1,2$. The set $\operatorname{Mod}_{\omega}(L_{i})$ carries a natural Borel structure, described in Section 2.5. In the formulation below we identify a code for a separable $L_{i}$-structure $A\in\operatorname{Mod}_{\omega}(L_{i})$ with the structure it encodes.
@@
Let $L$ be a countable metric language. Then the refining family of equivalence relations $(\equiv_{\alpha}^{bf})_{\alpha<\omega_{1}}$ on $\operatorname{Mod}_{\omega}(L)$ is uniformly Borel, and in particular uniformly analytic.
@@
To conclude Theorem A, we are left with showing that Corollary C can be applied to the set of classifiable $C^{*}$-algebras and the invariant $KT_{u}(\cdot)$. The descriptive set theory of the Elliott invariant, and of various assumptions of the unital classification theorem have been previously studied in [16]. In particular, using the results of Farah, Toms, and Törnquist, we reduce the proof of Theorem A to establishing that the set of $C^{*}$-algebras satisfying the UCT is an analytic set (with respect to a natural Borel structure on the set of unital separable $C^{*}$-algebras, see Section 3.1).
@@
Rosenberg and Schochet established their universal coefficient theorem (UCT) [49], which computes $KK$-theory [^3] For $C^{*}$-algebras $A,B$, $KK(A,B)$ is an abelian group that can be defined in terms of Kasparov $(A,B)$-bimodules. We use a different definition outlined in what follows. in terms of $K$-theory, for a large class of $C^{*}$-algebras. They show that a $C^{*}$-algebra satisfies the UCT if and only if it is $KK$-equivalent to an abelian $C^{*}$-algebra (for our purposes this can be taken as the definition of the UCT). This condition plays a central and mysterious role within the classification of $C^{*}$-algebras; it is unknown whether it is redundant and begs the open question of whether every simple separable nuclear $C^{*}$-algebra satisfies the UCT. One can think of $KK$-equivalence as a loose notion of homotopy equivalence between $C^{*}$-algebras, and we study the descriptive complexity of the UCT class by using the picture of $KK$ as a ‘universal functor’. More precisely, from the work [31] of Higson, it follows two $C^{*}$-algebras are $KK$-equivalent if and only if their images under any functor that satisfies certain conditions (that we describe in Section 3.1, see Definition 3.7) become equivalent. This allows us to present $KK$-equivalence of $C^{*}$-algebras $A,B$ in an existential way: there exist a finite number of conditions on functors $F$, that imply equivalence of images $F(A)\cong F(B)$.
@@
In order to show that $KK$-equivalence is an analytic relation (which implies analyticity of the UCT), we define a first-order language $L_{KK}$ and an $L_{KK}$-theory $T_{KK}$, whose models correspond to functors satisfying Higson’s conditions. We prove by a thorough analysis of Higson’s conditions that $T_{KK}$ is analytic as a subset of the set of $L_{KK}$-sentences with a suitable Borel space structure. This is carried out in Section 3.3 and uses a particular construction of the Borel category $\mathbf{C^{*}}$ of separable $C^{*}$-algebras. More precisely, in Section 2.5 we describe a Borel category equivalent to the category of separable models of a fixed metric $L_{\omega_{1},\omega}$-sentence with $L$-homomorphisms as morphisms, that we use in the case of $C^{*}$-algebras (see Proposition 2.39). Let us note that if one considers only surjective isometries between models as morphisms, such Borel groupoids were already considered in [12, 44] (using the Urysohn space) and in [2]. As a corollary of analyticity of $T_{KK}$, we get the following, which is a result of interest in its own right.
@@
    It follows from previous works on the descriptive set theory of $C^{*}$-algebra invariants (together with Theorem E) that the set of classifiable $C^{*}$-algebras is analytic, and that the map $KT_{u}(\cdot)$ taking a $C^{*}$-algebra to its invariant is Borel.
@@ Acknowledgements
We are indebted to Udi Hrushovski and Stuart White for copious amounts of guidance. We thank the participants of the reading group on [9] in Oxford, Michaelmas 2024, in particular Jakub Curda, Shanshan Hua and Austin Shiner. Moreover, we thank Filippo Calderoni, Leo Gitin, and Andrea Vaccaro for some conversations.
Funding. JP was supported by the Engineering and Physical Sciences Research Council (EP/X026647/1). MS was supported by the National Science Foundation under Grant No. DMS-2424441, and by the IAS School of Mathematics.
@@ 2. Descriptive set theory and infinitary metric logic
The model theory of metric structures, dating back to work of Chang and Keisler [6], was developed extensively in the past two decades, and large parts of (discrete) model theory have been generalized to the metric context [1, 4]. Metric model theory has since received growing attention, particularly within the study of operator algebras [14, 22, 21, 35, 48]. The study of infinitary logics emerged in the 1960s, with particular focus on the logic $L_{\omega_{1},\omega}$, where one allows disjunction and conjunction over countable sets of formulas. [^4] The subscript $\omega_{1}$ indicates that only countable disjunctions and conjunctions appear in $L_{\omega_{1},\omega}$-formulas, while the subscript $\omega$ indicates that only finite-length blocks of quantifiers appear.
We aim to give in Section 2.2 a brief introduction to the lesser studied area of infinitary metric logic as treated in [2], as well as its connection to descriptive set theory through considering spaces of separable models. We define the notions of a metric language $L$, the corresponding infinitary logic $L_{\omega_{1},\omega}$ and the quantifier rank $\operatorname{qr}(\varphi)$ of an $L_{\omega_{1},\omega}$-formula, and describe the standard Borel space $\operatorname{Mod}_{\omega}(L)$ of separable $L$-structures. We will assume familiarity with some basic concepts in discrete first-order logic.
In section 2.3 we give a presentation of the back-and-forth pseudo-distance of Ben Yaacov, Doucha, Nies, and Tsankov [2], in terms of a game which we refer to here as the dynamic metric Ehrenfeucht-Fraïssé game $\operatorname{EFD}_{\alpha}$. We show that the set of winning strategies for such games is Borel, and derive in particular that the family of equivalence relations $(\equiv_{\alpha}^{bf})_{\alpha<\omega_{1}}$ between separable structures is Borel in a uniform sense, see Theorem 2.21.